OpenStudy (anonymous):

A spinner is divided into sections shaded either black or gray. The probability that the spinner will stop on a black section is 20 percent. What is the probability that the spinner will stop on a gray section? A. 20% B. 40% C. 60% D. 80%

OpenStudy (anonymous):

kc_kennylau

OpenStudy (kc_kennylau):

1 - 20%

OpenStudy (kc_kennylau):

Because it's either black or gray

OpenStudy (kc_kennylau):

So P(black) + P(gray) = 1

OpenStudy (kc_kennylau):

Get it?

OpenStudy (anonymous):

no

OpenStudy (kc_kennylau):

P(something) means probability of something

OpenStudy (kc_kennylau):

P(A or B) = P(A) + P(B) P(A and B) = P(A) * P(B)

OpenStudy (kc_kennylau):

These two formulas will be very useful

OpenStudy (kc_kennylau):

It's either black or gray, so P(black or gray) = 1 because it must happen

OpenStudy (kc_kennylau):

Anything that "must happen" will have a probability of 1

OpenStudy (kc_kennylau):

Get it?

OpenStudy (kc_kennylau):

Do you understand why P(black or gray)=1

OpenStudy (anonymous):

yes

OpenStudy (kc_kennylau):

P(black or gray)=1 Therefore P(black) + P(gray) = 1 Do you understand this

OpenStudy (anonymous):

kinda

OpenStudy (anonymous):

those are complimentary events.

OpenStudy (kc_kennylau):

Because P(A or B) = P(A)+P(B)

OpenStudy (kc_kennylau):

Therefore P(black or gray) = P(black) + P(gray)

OpenStudy (kc_kennylau):

Therefore P(black) + P(gray) = 1

OpenStudy (kc_kennylau):

Now you have P(black)=20%

OpenStudy (kc_kennylau):

Substitute P(black)=20% to P(black)+P(gray)=1

OpenStudy (kc_kennylau):

Understand?

OpenStudy (anonymous):

i think

OpenStudy (kc_kennylau):

Then do it

OpenStudy (anonymous):

probability of wheel to stop at black is 20%. so what if doesnt stop on black .. it stops on grey .. it means its probability become rest of 100. that is 100-20 = 80

OpenStudy (anonymous):

ohhh now i understand

OpenStudy (kc_kennylau):

Glad you understand, because we gave no explanation to the questions before this

OpenStudy (anonymous):

so the ans id 80

OpenStudy (kc_kennylau):

Yes

OpenStudy (anonymous):

@kc_kennylau tried very hard to make you @72595210002 understand .. i give him my medal.

OpenStudy (kc_kennylau):

Thanks